Two islands A and B are 2.0km apart. How far is a third island C from A and from B is angle ACB=90 degrees



1
Expert's answer
2022-05-06T12:38:10-0400

"sin (\alpha)=AC/AB"

"AB=\sqrt{AC^2+BC^2}"

BC=AC

"2=\sqrt{2AC^2}"

"4=2AC^2"

"AC=BC=\sqrt2"